### v3

@@ 1 Study areas and disaster events
1.1 Explosion in Beirut, 2020
On August 4, 2020, a massive explosion occurred at the Port of Beirut in Lebanon, caused by the improper storage of 2,750 tons of ammonium nitrate. The explosion caused widespread damage within a radius of several kilometers, significantly impacting the port and the surrounding neighborhoods, including areas such as Gemmayzeh, Mar Mikhael, and Achrafieh. It resulted in 218 deaths, more than 7,000 injuries, and left approximately 300,000 people homeless (Fakih and Majzoub, 2021). Economic losses were estimated to be 15 billion USD (Valsamos et al., 2021). The disaster compounded Lebanon’s ongoing economic challenges and contributed to political instability and social unrest.
1.2 Explosion in Bata, 2021
On March 7, 2021, a series of four explosions occurred at the Cuartel Militar de Nkoantoma in Bata, Equatorial Guinea, caused by improperly stored explosives. The blasts led to at least 107 deaths, over 615 injuries, and widespread destruction throughout the city (OCHA, 2021a). A total of 243 structures were destroyed or severely damaged, displacing many residents. Around 150 families sought refuge in temporary shelters, while others stayed with relatives. Local hospitals treated more than 500 injured individuals, and the economic impact was severe, underscoring the dangers associated with the improper storage of hazardous materials.
1.3 Volcano Eruption in DR Congo and Rwanda, 2021
On May 22, 2021, Mount Nyiragongo in the Democratic Republic of the Congo erupted, causing widespread devastation. The eruption resulted in 32 deaths, the destruction of 1,000 homes, and the displacement of thousands as lava flows threatened the city of Goma (IFRC, 2021). Nearly 400,000 people were evacuated due to the risk of further volcanic activity, including potential magma flow beneath Goma and nearby Lake Kivu. Despite continued seismic activity, life in Goma largely returned to normal by August 2021, although plans to relocate parts of the city remain under consideration due to the ongoing threat from the volcano.
1.4 Earthquake in Haiti, 2021
On August 14, 2021, a magnitude 7.2 earthquake struck Haiti’s Tiburon Peninsula, primarily affecting the Nippes, Sud, and Grand’Anse departments. The disaster caused over 2,200 deaths, more than 12,200 injuries, and left thousands homeless. The economic losses were significant, estimated at over USD 1.5 billion. Approximately 137,500 buildings, including homes, schools, and hospitals, were damaged or destroyed (OCHA, 2021b). As the deadliest natural disaster of 2021, the earthquake exacerbated Haiti’s existing challenges, including widespread poverty and political instability.
1.5 Volcano Eruption in La Palma, 2021
On September 19, 2021, the Cumbre Vieja volcano on La Palma, part of Spain’s Canary Islands, erupted following several days of seismic activity. The eruption primarily impacted the island’s western side, covering over 1,000 hectares with lava and destroying more than 3,000 buildings, including the towns of Todoque and La Laguna. The lava flow, measuring about 3.5 kilometers wide and 6.2 kilometers long, reached the sea, cutting off the coastal highway and forming a new peninsula with extensive lava tubes. Although the timely evacuation of around 8,000 people prevented major casualties, one person died from inhaling toxic gases (Troll et al., 2024). Economic losses exceeded 800 million Euros, and thousands of residents were displaced, with significant damage to arable land and livelihoods.
1.6 Wildfire in Colorado, 2021
The Marshall Fire, which started on December 30, 2021, in Boulder County, Colorado, became the most destructive wildfire in the state history in terms of destroyed buildings. Fueled by dry grass from an unusually warm and dry season, and winds up to 185 km/h, the fire killed two people and injured eight. It destroyed 1,084 structures, including homes, a hotel, and a shopping center, causing over US\$2 billion in damage (Forrister et al., 2024). More than 37,500 residents evacuated, and significant damage to public drinking water systems occurred (Forrister et al., 2024).
1.7 Armed Conflict in Ukraine, 2022
In February 2022, Russian forces launched a full-scale invasion of Ukraine, resulting in widespread destruction and displacement. By November 2024, total damages to Ukraine’s infrastructure had reached \$170 billion. More than 236,000 residential buildings were damaged or destroyed, including 209,000 private houses, 27,000 apartment buildings, and 600 dormitories. Over 4,000 educational institutions and 1,554 medical facilities were also affected, alongside extensive damage to transport, energy, and telecommunications infrastructure. The conflict has inflicted more than 40,000 civilian casualties, displaced four million people internally, and forced 6.8 million refugees to flee. As of late 2024, approximately 14.6 million Ukrainians require humanitarian assistance (Andrienko et al., 2025).
1.8 Earthquake in Turkey, 2023
On February 6, 2023, a magnitude 7.8 earthquake struck southeastern Turkey near Gaziantep, followed by a magnitude 7.7 aftershock. The disaster, the most powerful earthquake in Turkey since 1939, caused widespread destruction across approximately 350,000 $km^{2}$, affecting 14 million people and displacing 1.5 million. The death toll reached 53,537, with 107,213 injuries (STL, 2024), making it one of the deadliest earthquakes in modern history. Economic losses were estimated at USD 148.8 billion, with over 518,000 houses and 345,000 apartments destroyed (Government of Türkiye, 2023). The earthquake caused severe damage to infrastructure, agriculture, and essential services, further worsening the region’s economic challenges. International aid was mobilized to support the affected populations.
1.9 Cyclone in Myanmar, 2023
In May 2023, Cyclone Mocha, a Category 5 hurricane, struck Myanmar, causing widespread devastation in the country’s coastal regions, particularly in Rakhine State. According to official estimates, at least 148 people were killed and 132 injured, although other sources suggest higher figures (GFDRR and World Bank, 2023). The cyclone affected around 1.2 million people in Rakhine alone, with over 200,000 buildings reportedly damaged or destroyed, making it one of the most destructive cyclones in the region in the past 15 years. The direct economic damage was estimated at US\$2.24 billion in damages, equivalent to 3.4% of Myanmar’s GDP (GFDRR and World Bank, 2023).
1.10 Wildfire in Hawaii, 2023
In August 2023, a series of wildfires broke out on the island of Maui, Hawaii, causing widespread destruction and significant impacts on the local population and environment. The fires, fueled by dry conditions and strong winds, primarily affected the town of Lahaina, where at least 102 people were killed and two remain missing (Hedayati et al., 2024). Over 2,200 buildings were destroyed, including many historic landmarks, resulting in estimated damages of US\$5.5 billion (Jones et al., 2024; NCEI, 2025). The fires prompted evacuations and led to the displacement of thousands of residents, with significant economic losses in the tourism and agriculture sectors.
1.11 Earthquake in Morocco, 2023
On September 8, 2023, a 6.9 magnitude earthquake struck Morocco’s Al Haouz Province near Marrakesh, causing widespread devastation. The earthquake killed nearly 3,000 people, injured more than 5,500 and displaced more than half a million (OCHA, 2023). The earthquake damaged or destroyed nearly 60,000 houses, with the heaviest losses reported in rural communities of the Atlas Mountains (International Federation of Red Cross and Red Crescent Societies, 2024). Overall damage is estimated at about USD 7 billion (National Geophysical Data Center / World Data Service, 2025), while direct economic losses of roughly USD 30 million amount to approximately $0.24\%$ of Morocco’s GDP. Cultural heritage in Marrakesh was also hard-hit: sections of the UNESCO-listed Medina and several historic mosques sustained severe damage.
1.12 Flood and Storm in Libya, 2023
In September 2023, Storm Daniel brought catastrophic flooding to northeastern Libya, particularly in Derna, when two dams collapsed and released an estimated 30 million cubic meters of water. At least 5,923 people were killed, though local officials warned fatalities could reach 18,000–20,000 (USAID Bureau for Humanitarian Assistance, 2024). Post–event analyses revealed that approximately $10\%$ of Derna’s housing stock was destroyed and a further $18.5\%$ sustained damage. Across the wider coastal belt—from Benghazi, through Jabal Al Akhdar and Al Marj, to Derna—an estimated 18 838 dwellings were damaged or obliterated (Normand and Heggy, 2024).. The disaster, considered the second deadliest dam failure in history, devastated the city’s infrastructure: four bridges collapsed, and entire neighborhoods were washed out to sea. Long-standing neglect of dam maintenance, compounded by Libya’s political turmoil, contributed significantly to the scale of the tragedy.
1.13 Hurricane in Mexico, 2023
In September 2023, Hurricane Norma, a Category 4 hurricane, struck the western coast of Mexico, severely affecting Sinaloa and Baja California Sur. This was followed by Hurricane Otis in October, which made landfall near Acapulco as a Category 5 hurricane. Otis was the strongest Pacific hurricane to hit Mexico, causing at least 52 deaths and leaving 32 missing. The storm caused unprecedented destruction, with more than 51,864 homes destroyed and damages estimated at US\$12–16 billion, surpassing Hurricane Wilma as the costliest Mexican hurricane (Reinhart and Reinhart, 2024).
1.14 Earthquake in Noto, 2024
On January 1, 2024, a magnitude 7.5 earthquake struck the Noto Peninsula in Ishikawa Prefecture, Japan, reaching a maximum JMA seismic intensity of Shindo 7. The earthquake caused widespread destruction, particularly in the towns of Suzu, Wajima, Noto, and Anamizu, and triggered a 7.45-meter tsunami along the Sea of Japan coast. The disaster resulted in 572 deaths and over 1,300 injuries, with 193,529 structures damaged across nine prefectures (Fire and Disaster Management Agency, 2025). It was the deadliest earthquake in Japan since the 2011 Tōhoku disaster. As of February 20, 12,929 people remained in 521 evacuation centers (Japanese Red Cross Society, 2024). The event prompted Japan’s first major tsunami warning since 2011. Total damage is estimated 16.9 trillion yen, reflecting substantial economic losses with significant impacts on infrastructure and communities across the affected regions (Kaneko, 2024).
@@ 1.14 Earthquake in Noto, 2024
Figure 4: Thumbnails of local areas in 14 disaster events of Bright. For visualization purposes, different events have different scales.
@@ 2 Dataset Description
2.1 Construction of the dataset
Figure 3 shows the flowchart of developing Bright. The optical EO data in our dataset are mainly from Maxar’s Open Data program [^3] https://www.maxar.com/open-data, while the SAR EO data are from Capella Space [^4] https://www.capellaspace.com/earth-observation/gallery and Umbra [^5] https://umbra.space/open-data/. Both Capella and Umbra data have two imaging modalities, i.e., Spotlight and Stripmap, respectively. The Spotlight mode has a higher spatial resolution but less coverage. In the region of interest, we preferred Spotlight mode if there is suitable data in the data provider’s inventory. Otherwise, we chose Stripmap. The optical EO data consists of red, blue, and green bands, while the SAR EO data are amplitude data in the VV or HH bands. For optical EO data, the digital number was converted to reflectance and then standardized to an 8-bit data format. For SAR imagery, after the data had been terrain-corrected, we used the data provider’s recommended method of pre-processing the amplitude data and finally converted the data to an 8-bit data format. Although both optical and SAR images are geocoded, there are still pixel offsets between them. Therefore, multiple EO experts manually aligned the paired optical and SAR data and cross-checked their results to ensure the precise registration between the two types of EO data. Figure 11 shows the selected control points on three disaster scenes.
Table 3: Definition of the different categories of damage in Bright.
@@ 2.1 Construction of the dataset
The labels in Bright consist of two components: building polygons and post-disaster building damage attributes. Expert annotators manually labeled the building polygons, and all labels underwent independent visual inspections of EO experts to ensure accuracy. Damage annotations were obtained from Copernicus Emergency Management Service [^6] https://emergency.copernicus.eu, the United Nations Satellite Centre (UNOSAT) Emergency Mapping Products [^7] https://unosat.org/products, and The Federal Emergency Management Agency (FEMA) [^8] https://www.fema.gov. These annotations were derived through visual interpretation of high-resolution optical imagery captured before and after the disasters by EO experts, supplemented by partial field visits. To further improve damage information reliability, we performed a secondary validation and correction of these annotations. EO experts carefully reviewed the damage points using multi-temporal imagery available on platforms such as Google Earth and manually refined the labels when inconsistencies or misclassifications were observed. Since the number of damage levels varied across different events, we refered to the framework in FEMA’s Damage Assessment Operations Manual, EMS-98, the BDD dataset (Adriano et al., 2021), and the xBD dataset (Gupta et al., 2019) to standardize them into three categories: intact (1), damaged (2), and destroyed (3). Table 3 summarizes these definitions. The damage annotations were provided as vector point files. By overlaying these points with the building polygons and assigning corresponding damage attributes, we generated the final building damage labels. To prevent geographic misallocations due to possible coordinate offsets, we unified the coordinate systems of the points and polygons, with a visual inspection performed prior to the final allocation.
Figure 4 presents thumbnails of selected local areas from the 14 disaster events.
2.2 Statistics for Bright
@@
Figure 5: Statistics for the Bright dataset. (a) Distributions of band values of samples from four study sites. (b) Distribution of building scales. (c) Feature distribution of buildings on four events under two imaging modalities. (d) Percentage of building and background pixels and percentage of different damage levels in building pixels.
@@
The basic information about Bright, including disaster events, EO data, the number of corresponding EO tiles, and the total number of building pixels, is summarized in Table 2. After cropping the EO data into 1024$\times$1024-pixel tiles, Bright contains 4,246 multimodal image pairs.
@@
The key statistics of Bright are illustrated in Figure 5. Figure 5a shows the pixel value distribution for optical and SAR images from one human-made disaster and three natural disasters. The varying geographical landscapes and land cover across different regions result in distinct means and standard deviations of pixel values, highlighting Bright’s geographical diversity and making it a robust dataset for studying building damage assessment in diverse environments. To ensure trained models can accurately detect buildings and assess damage levels, it is crucial that the dataset includes a wide variety of building styles from different regions. Figure 5b shows that Bright covers buildings at multiple scales, exhibiting a “long-tail” distribution. This multi-scale representation challenges DL models to develop the ability to capture features at varying scales, enhancing robustness and accuracy.
@@
Table 4: Proxy registration errors (in pixel) estimated using different multimodal image registration methods. The estimated registration errors for each event are reported in Table 13.
@@
We also evaluated the registration accuracy between optical and SAR EO data in Bright. Due to the absence of real ground truth, we introduce a proxy method to estimate registration errors, leveraging existing multimodal image descriptors. This approach is detailed in Appendix B. Table 4 reports the mean registration errors, measured as the root mean square error (RMSE) of pixel displacements, obtained using three representative multimodal image registration methods: RIFT (Li et al., 2020), SRIF (Li et al., 2023a), and LNIFT (Li et al., 2022). The overall average RMSE is approximately 1.024 pixels, with a lower error of 1.006 pixels specifically within building regions. Accurate registration results ensure spatial consistency across modalities, providing a solid foundation for accurate building damage assessment.
2.3 Dataset division strategy
@@
To train DL models using Bright and evaluate their generalizability, it is necessary to divide the dataset into a training set, validation set, and test set. Gerard et al. (2024) suggested that dividing the dataset on an event-by-event basis, rather than randomly across the entire dataset, provides a more accurate reflection of a model’s generalizability. Therefore, for the 14 events listed in Table 2, we divide the corresponding data for each event into training, validation, and test sets in a ratio of 7:1:2. The sets obtained for each event are then merged to create the final training, validation, and test sets. In our experiments, DL models are trained using the training set, and the optimal hyperparameters (e.g., learning rate) and checkpoints are selected based on performance on the validation set. The generalization capability of the models is subsequently evaluated on the test set.
@@
In addition to the above standard machine learning division, we also introduce a cross-event transfer setup to better evaluate the ability of models to generalize across disaster events, a critical challenge in real-world applications where models are expected to handle unseen disaster types and locations. This setup comprises two key divisions:
@@
  • Zero-shot setup: This setting mimics a real-world scenario where a newly occurring disaster must be analyzed without any prior labeled data from the same event. We isolate one event as an unseen test set while using the remaining 13 events for training and validation. This setting evaluates the cross-event generalization ability of models, testing how well they can transfer learned knowledge from previous disasters to an entirely new event. Due to the high variability of disaster types and geographies, this setup is inherently challenging, as models trained on past disasters may struggle to assess damage patterns in a previously unseen event accurately.
@@
  • One-shot setup: Recognizing the difficulty of the zero-shot setup, we introduce a one-shot setup, acknowledging that in real-world disaster response, a small number of labeled samples from the new disaster can be obtained. In this setting, a limited subset of labeled data (one pair for training and one pair for validation) from the target disaster event is incorporated into the training process. At the same time, the majority of the test set remains unseen. This setup evaluates the model’s ability to leverage a minimal amount of manually labeled data to improve disaster-specific adaptation.

### v4

@@ 0.2 Main contribution
Figure 4: Thumbnails of local areas in 14 disaster events in the Bright dataset. The sources of EO images are illustrated in Table 2. For visualization purposes, different events have different scales.
@@ 1 Dataset Description
1.1 Study areas and disaster events
We selected 14 disaster events across the globe for Bright, as illustrated in Figure 2 and Table 2. Since both Capella Space and Umbra satellites were launched in 2020, we focused on study areas where disasters have occurred since then. The selected regions are primarily in developing countries, where public administration and disaster response capacities tend to be weaker compared to those in developed nations, making international assistance more critical. The dataset covers five major types of natural disasters: earthquakes, storms (including hurricanes and cyclones), wildfires, floods, and volcanic eruptions. Additionally, it includes human-made disasters, such as accidental explosions and armed conflicts. Detailed descriptions of the 14 disaster events are provided in Appendix A.
1.2 Construction of Bright
Figure 3 shows the flowchart of developing Bright. The optical EO data in the dataset are mainly from Maxar’s Open Data program [^3] https://www.maxar.com/open-data, while the SAR EO data are from Capella Space [^4] https://www.capellaspace.com/earth-observation/gallery and Umbra [^5] https://umbra.space/open-data/. Both Capella and Umbra data have two imaging modalities, i.e., Spotlight and Stripmap, respectively. The Spotlight mode has a higher spatial resolution but less coverage. In the region of interest, we preferred Spotlight mode if suitable data was available in the data provider’s inventory. Otherwise, we chose Stripmap. The optical EO data consists of red, blue, and green bands, while the SAR EO data consists of amplitude data in the VV or HH bands. For optical EO data, the digital number was converted to reflectance and then standardized to an 8-bit data format. For SAR imagery, after the data had been terrain-corrected, we utilized the pre-processed 8-bit data when available. In cases where 8-bit data was not provided, we employed the data provider’s recommended method [^6] https://support.capellaspace.com/scaling-geo-images-in-qgis to convert the amplitude data. Although both optical and SAR images are geocoded, there are still pixel offsets between them. Therefore, multiple EO experts manually aligned the paired optical and SAR data and cross-checked their results to ensure the precise registration between the two modalities. Figure 12 in Appendix C shows the selected control points on three disaster scenes.
Table 3: Definition of different categories of damage in Bright.
@@ 1.2 Construction of Bright
The labels in Bright consist of two components: building polygons and post-disaster building damage attributes. Expert annotators manually labeled the building polygons, then all labels underwent independent visual inspections of EO experts to ensure accuracy. Damage annotations were obtained from Copernicus Emergency Management Service [^7] https://emergency.copernicus.eu, the United Nations Satellite Centre (UNOSAT) Emergency Mapping Products [^8] https://unosat.org/products, and the Federal Emergency Management Agency (FEMA) [^9] https://www.fema.gov. These annotations were derived through visual interpretation of high-resolution optical imagery captured before and after the disasters by EO experts, supplemented by partial field visits. To harmonize these diverse annotations and ensure consistency across all 14 disaster events, we implemented a rigorous, multi-stage process. First, we established a single, standardized three-tier classification scheme, including Intact (with pixel value 1), Damaged (with pixel value 2), and Destroyed (with pixel value 3), with clear definitions provided in Table 3, drawing on the frameworks of FEMA’s Damage Assessment Operations Manual, EMS-98, the BDD dataset (Adriano et al., 2021), and the xBD dataset (Gupta et al., 2019). While the source agencies’ terminology can differ (e.g., “Severe Damage” vs. “Major Damage”), their underlying definitions for EO-based assessment are conceptually consistent. We leveraged this alignment for an initial rule-based mapping, where various intermediate damage tiers were conservatively aggregated into our single “Damaged” category. Second, our team of EO experts conducted a comprehensive manual verification and refinement of every annotation using multi-temporal VHR imagery on platforms like Google Earth Pro. This final stage served as the ultimate guarantor of consistency. We paid special attention to ambiguous source labels, such as “Possibly Damaged”. Adopting a conservative approach, these were re-classified as “Intact” if clear structural damage was not evident, thereby ensuring a high-confidence “Damaged” class. We also manually disaggregated all area-based annotations (i.e., where an entire block was assigned a single category). We re-processed these to assign a precise, building-wise damage label to each individual structure, ensuring instance-level consistency and granularity across the entire dataset. The damage annotations were provided as vector point files. The final building damage labels were generated by overlaying these points withd the building polygons and assigning corresponding damage attributes. To prevent geographic misallocation due to possible coordinate offsets, the coordinate systems of the points and polygons were unified, with a visual inspection performed prior to the final allocation. Figure 4 presents thumbnails of selected local areas from the 14 disaster events.
1.3 Statistics of Bright
@@
Figure 5: Statistics of the Bright dataset. (a) Distributions of band values of samples from four study sites. (b) Distribution of building scales. (c) Feature distribution of buildings of four events under two imaging modalities. (d) Percentage of building and background pixels and percentage of different damage levels in building pixels.
@@
The basic information about Bright, including disaster events, EO data, the number of corresponding EO tiles, and the total number of building pixels, is summarized in Table 2. After cropping the EO data into $\times1024$-pixel tiles, Bright contains 4,246 multimodal image pairs.
@@
The key statistics of Bright are illustrated in Figure 5. Figure 5-(a) shows the pixel value distribution for optical and SAR images from one human-made disaster and three natural disasters. The varying geographical landscapes and land cover across different regions result in distinct means and standard deviations of pixel values. This highlights Bright’s geographical diversity, which makes it a robust dataset for studying building damage assessment in diverse environments. To ensure that models trained on Bright can accurately detect buildings and assess damage levels, it is crucial that the dataset includes a wide variety of building styles from different regions. Figure 5-(b) shows that Bright covers buildings at multiple scales, exhibiting a “long-tail” distribution. This multi-scale representation challenges DL models to develop the ability to capture features at varying scales, enhancing robustness and accuracy.
@@
Table 4: Proxy registration errors (in pixel) estimated using different multimodal image descriptors. The estimated registration errors for each event are reported in Table 15 in Appendix C.
@@
Moreover, since accurate registration ensures spatial consistency across modalities, the registration accuracy between optical and SAR EO data in Bright was analyzed to provide a solid foundation for accurate building damage assessment. Due to the absence of real ground truth, a proxy method that leverages existing multimodal image descriptors was introduced to estimate registration errors. This approach is detailed in Appendix C. The Table 4 reports the mean registration errors, measured as the root mean square error (RMSE) of pixel displacements. It was obtained using three representative multimodal image registration methods: RIFT (Li et al., 2020), SRIF (Li et al., 2023a), and LNIFT (Li et al., 2022). The overall average RMSE is approximately 1.024 pixels, with a lower error of 1.006 pixels specifically within the building regions.
1.4 Dataset splitting strategy
@@
To train DL models using Bright and evaluate their generalizability, it is necessary to split the dataset into a training set, validation set, and test set. Gerard et al. (2024) suggested that dividing the dataset on an event-by-event basis, rather than randomly across the entire dataset, provides a more accurate reflection of a model’s generalizability. Therefore, for the 14 events listed in Table 2, we divide the corresponding data for each event into a ratio of 7:1:2 for training, validation, and test subsets, respectively. Then, the subsets obtained for each event are merged to create the final training, validation, and test sets. In the experiments, the baseline models are trained using the training set, and the optimal hyperparameters (e.g., learning rate) and checkpoints are selected based on performance on the validation set. The generalization capability of the baseline models is subsequently evaluated on the test set.
@@
In addition to the above standard ML data splitting, we also introduce a cross-event transfer setup to better evaluate the ability of models to generalize across disaster events. This is a critical challenge in real-world applications where models are expected to handle unseen disaster types and locations. Two setups are established for cross-event transfer generalization:
@@
  • Zero-shot setup: This setting mimics a real-world scenario where a newly occurring disaster must be analyzed without any prior labeled data from the same event. We isolate one event as an unseen test set while using the remaining 13 events for training and validation. This setting evaluates the cross-event generalization ability of models, testing how well learned knowledge can be transferred from previous disasters to an entirely new disaster event. Due to the high variability of disaster types and geographies, this setup is inherently challenging, as models trained on past disasters may struggle to assess damage patterns in a previously unseen event accurately.
@@
  • One-shot setup: Recognizing the difficulty of the zero-shot setup, we introduce a one-shot setup. This setting simulates a realistic scenario where a single, representative sample from the new disaster can be quickly labeled to guide model adaptation. In this setting, a limited subset of labeled data (one pair for training and one pair for validation) from the target disaster event is incorporated into the training process. At the same time, the majority of the test set remains unseen. This setup evaluates the model’s ability to leverage a minimal amount of manually labeled data to improve disaster-specific adaptation.
It is worth noting that our cross-event transfer setup differs from classic few-shot learning tasks in the computer vision field (Shaban et al., 2017; Wang et al., 2020). Our goal is not to recognize new classes, but to adapt the model’s knowledge of existing classes to a new domain, i.e., an unseen disaster event.